Please help me solve for x in the following problem.
2(3x-1) -4(2-x) = 3x+1

Answer by unlockmath(1688)

Hello, Let's first expand 2(3x-1) -4(2-x) = 3x+1 this out to be:
6x-2-8+4x=3x+1 This can be rewritten as:
10x-10=3x+1 Subtract 3x from both sides and add 10 to both sides to give us:
7x=11 Divide 7 into both sides will be:
x=11/7
Make sense?
You can check it out by plugging the answer into the original equation.
RJ
Check out a new book I wrote at:
www.math-unlock.com